Sulfur (1400 kg) and pearl caustic soda (950 kg) are reacted to form sodium polysulphide (Na2Sx) to a total of 4000 L (balanced by H2O). What percentages (%) of total sulphur (S), sodium sulphide (Na2S), sodium thiosulphate (Na2S2O3) and sodium sulfur compounds are in solution?


1
Expert's answer
2021-02-26T05:55:41-0500

"4S+6NaOH\rightarrow Na_2S_2O_3+2Na_2S+3H_2O"


Moles of Sulphur = "\dfrac{1400}{32}=43.75 kmol"

Moles of NaOH = "\dfrac{950}{40}=23.75kmol"

From reaction,

6 moles of NaOH react with 4 moles of S

23.75 kmol of NaOH reacts with 15.833 kmol of S

Moles of S left = 43.75 - 15.833 = 27.91 kmol

Total volume of Solution = 4000 L

Moles of Na2S2O3 formed = 3.95 kmol

Moles of Na2S formed = 7.91 kmol

Moles of H2O formed = 11.87 kmol

Mass of S left = "27.91\times32=893.12kg"

Mass of Na2S2O3 formed = "3.95\times158=624.1kg"

Mass of Na2S formed = "7.91\times78=616.98 kg"

Mass of H2O formed = "213.66kg"

Mass/Volume percentage of components,

"S=\dfrac{893.12}{4000}\times100=22.328\%"

"Na_2S_2O_3=\dfrac{624.1}{4000}\times 100=15.6\%"

"Na_2S=\dfrac{616.98}{4000}\times100=15.42\%"

"H_2O=100-(22.328+15.6+15.42)=46.652\%" (since H2O is added externally also)